# Bipartite OTOC in open quantum systems: information scrambling and irreversibility

### v1

@@ I Introduction
The relationship between the classical and quantum dynamics of chaotic systems is at the heart of the study of quantum chaos. Recently, there has been a resurgence of interest in many-body quantum chaos and its relationship to information scrambling. Information scrambling refers to the spreading of quantum information over a complex quantum system such that it is not accessible to simple probes. Interest in this field was rejuvenated recently by works on black hole information paradox [1, 2], and the problem of quantum thermalization [3]. Traditionally, the ideas of random matrix theory were employed to probe quantum chaos [4, 5]. In recent times, the out-of-time-ordered correlator (OTOC) has emerged as a prominent tool to investigate information scrambling. This was first brought out in the context of superconductivity [6], and was popularized by [7]. The growth rate of OTOC is commonly attributed to the classical Lyapunov exponent [8]. OTOC has been extensively used in the study of a number of systems, such as in quantum field theories [9], random unitary models [10], spin chains [11, 12], and quantum optical models [8, 13, 14], among others [15, 16, 17]. Various forms of OTOC, such as regularized and physical OTOC, were also used in the context of the fluctuation-dissipation theorem, generalized to the case of OTOC [18].
@@ III.1 Dicke model
$\frac{\partial}{\partial t}A=$ $\mathcal{L}(A)=i[H_{Dicke},A]+\gamma(N_{thA}+1)\big{[}J_{+}AJ_{-}-\frac{1}{2}\{A,J_{+}J_{-}\}\big{]}$
$+\gamma N_{thA}\big{[}J_{-}AJ_{+}-\frac{1}{2}\{A,J_{-}J_{+}\}\big{]}$
$+\kappa(N_{thB}+1)\big{[}a^{\dagger}Aa-\frac{1}{2}\{A,a^{\dagger}a\}\big{]}$
$+\kappa N_{thB}\big{[}aAa^{\dagger}-\frac{1}{2}\{A,a^{\dagger}a\}\big{]},$ (12)
@@
where $\gamma$ acts as a dissipating factor modelling spontaneous emission, and $\kappa$ is responsible for cavity decay. Further, $N_{th_{k}}=\frac{1}{e^{\beta\omega_{k}}-1}$ with $\beta=\frac{1}{k_{B}T}$, and $k=A$, $B$ denoting the subspaces $A$ and $B$, respectively. By vectorizing this master equation, the Lindbladian superoperator $\mathcal{L}$ and, in turn, the CPTP map $\mathcal{E}^{\dagger}=e^{t\mathcal{L}}$ is obtained. We consider the atomic system to belong to the $A$ subspace, and the radiation field is relegated to the $B$ subspace. By applying the swap operators for the specific dimensions, the bipartite OTOC is obtained. Further, the Fock space of the field part is truncated in such a manner that increasing the dimension of the Fock space after a point doesn’t affect the dynamics  [13, 62]. The dynamics of the bipartite OTOC for the Dicke model for different values of atom-field interaction strength $\lambda$ is shown in Fig. 2.
@@
Figure 2: Variation of averaged bipartite OTOC for Dicke model $G(\mathcal{E}^{\dagger})$ with time $t$ for different values of spin-cavity field interaction strength $\lambda$. The value of parameters are: $\omega_{0}=2$, $\omega_{c}=2$, $T=2$, $\gamma=\kappa=0.05$.
@@
We observe that as the atom-field coupling strength $\lambda$ increases, the fluctuations (or revivals) in the Bipartite OTOC $G(\mathcal{E}^{\dagger})$ decrease; that is, the curves smooth out. This is predominant when the coupling parameter approaches the critical value $\lambda_{c}=\frac{\sqrt{\omega\omega_{0}}}{2}$, where the Dicke model transitions from the normal regime to the superradiant regime. In [13], an exponential behavior of OTOC was observed when the Dicke model underwent a phase transition from the normal to the superradiant regime. A consistent result is observed here for the bipartite OTOC, where the smoothening of the curve in the superradiant regime, characterized here by $\lambda=1$, could be considered an indicator of the onset of quantum chaos. Further, in Fig. 3, we plot the variation of the bipartite OTOC for different temperatures for a given value of $\lambda$. It is observed that the bipartite OTOC takes the highest values for zero temperature. As the temperature increases, the values of the bipartite OTOC reduce.
Figure 3: Bipartite OTOC for Dicke model with coupling strength $\lambda=1.5$ for four different temperatures. The value of parameters are: $\omega_{0}=2$, $\omega_{c}=2$, $\gamma=\kappa=0.05$.
@@ III.2 Tavis-Cummings model
The total number of the excitations $\mathcal{N}=a^{\dagger}a+J_{z}$ in this model remains conserved. The sources of dissipation in this model are the same as in the Dicke model. By replacing $H_{Dicke}$ in Eq. (III.1) by $H_{TC}$, we arrive at an adjoint master equation for the TC model. The swap operators for this model remain the same as those for the Dicke model.
Figure 4: Variation of averaged bipartite OTOC for Tavis Cummings model $G(\mathcal{E}^{\dagger})$ with time $t$ for different values of spin-cavity field interaction strength $\lambda$. The value of parameters are: $\omega_{0}=2$, $\omega_{c}=2$, $T=2$, $\gamma=\kappa=0.05$.
Figure 4 depicts the variation of the bipartite OTOC for the TC model. It can be observed that, like the Dicke model, as the coupling strength increases, the fluctuations (or revivals) in the bipartite OTOC $G(\mathcal{E}^{\dagger})$ decrease. The fluctuations in the bipartite OTOC act as an identifier of an integrable system; the smoothness of the bipartite OTOC indicates non-integrability. The least revivals are seen when the coupling parameter is around the critical value $\lambda_{c}=\sqrt{\omega\omega_{0}}$ (2 for the parameters chosen here) for a phase transition from the normal to the superradiant regime. This is again consistent with what was observed in [13].
@@ III.3 Ising model with tilted magnetic field
$H_{A}$ $=\mathcal{B}(sin(\theta)\sigma^{x}_{1}+cos(\theta)\sigma^{z}_{1}),$
$H_{B}$ $=\mathcal{B}\sum_{i=2}^{N}(sin(\theta)\sigma^{x}_{i}+cos(\theta)\sigma^{z}_{i}),~{}~{}~{}~{}\text{and}$
@@
$\frac{\partial}{\partial t}A$ $=i[H_{\theta},A]+\sum_{j=1}^{N}\bigg{(}\gamma(N_{th_{j}}+1)\bigg{[}\sigma_{+}^{j}A\sigma_{-}^{j}-\frac{1}{2}\{A,\sigma_{+}^{j}\sigma_{-}^{j}\}\bigg{]}$
$+\gamma N_{th_{j}}\bigg{[}\sigma_{-}^{j}A\sigma_{+}^{j}-\frac{1}{2}\{A,\sigma_{-}^{j}\sigma_{+}^{j}\}\bigg{]}\bigg{)},$ (16)
where $N$ is the number of spins in the Ising chain being acted upon by the bath, $\gamma$ is the dissipation factor, and $N_{th_{j}}=\frac{1}{e^{\beta\omega_{j}}-1}$.
@@
The variation of the bipartite OTOC for this model for four different angles of tilt, including the two extreme cases of $\theta=0^{\circ}$ and $\theta=90^{\circ}$ is plotted in Fig. 6.
Figure 6: Variation of averaged bipartite OTOC for the Ising model in tilted magnetic field $G(\mathcal{E}^{\dagger})$ with time $t$ for different values of angle of tilt of the magnetic field $\theta$. The value of parameters are: $\omega_{0}=2$, $\omega_{c}=2$, $T=2$, $\gamma=0.05,\mathcal{B}=J=0.5$.
@@
Further, we study the bipartite OTOC in the above Ising model under the assumption that the bath acts only on the first and the last spins of the Ising chain. The bipartition of the system is taken as 2 spins in the $A$ subsystem and 3 spins in the $B$ subsystem, as illustrated in the second sketch of Fig. 5. Here, we can study a spin chain attached to the baths at the same or different temperatures on either side. In this case, only the first and last spins are subjected to the dissipating Lindblad operators $L_{j}$’s as $\sigma_{\pm}^{j}$ (where $j=1,N)$. To this end, the above master equation [Eq. (III.3)] with appropriate Lindblad operators is used. The bipartite OTOC for this system is shown in Fig. 7 for two different angles of tilt with the magnetic field for the same or different temperatures of the two baths.
Figure 7: Bipartite OTOC for a system with two bipartitions with subsystem A having 2 spins and B having 3 spins with the first and last spins experiencing the influence of baths. Two cases are considered where the two baths are at the same ($T_{1}=T_{2}=1$) or different temperatures ($T_{1}=1$, and $T_{2}=10$). Further, $J=\mathcal{B}=1$.
It is evident from the figure that the bipartite OTOC takes lesser values in the presence of different bath temperatures, with the last spin having a higher temperature than the first one. Furthermore, a smooth graph of the bipartite OTOC is obtained for the angle $167\pi/$ 16, indicating non-integrable behavior, whereas the graph for the angle $\pi/2$ has fluctuations in it, indicative of integrability. This is consistent with the (non-)integrable behavior of the Ising model, Eq. (III.3).
@@ IV Bipartite OTOC, operator entanglement, and entropy production
The operator entanglement in many-body quantum systems quantifies how much the local operators become entangled over time. It serves as a natural candidate for distinguishing between chaotic and integrable dynamics in quantum systems [64]. Further, it was shown in [19, 22, 21] that the bipartite OTOC for a unitary evolution $U$ matches the operator entanglement $E_{op}(U)$. The total unitary |$\ket{U}_{ABA^{\prime}B^{\prime}}$ is supported over the composite Hilbert space $\mathcal{H}_{AB}$ and its replica $\mathcal{H}_{A^{\prime}B^{\prime}}$. Since the entanglement of state |$\ket{U}_{ABA^{\prime}B^{\prime}}$ across $AA^{\prime}|BB^{\prime}$ is non-trivial, in contrast to the bipartition $AB|A^{\prime}B^{\prime}$ which is maximal due to it being local unitarily equivalent to the maximally entangled state, $B$ and $B^{\prime}$ are traced out. The explicit expression for $E_{op}(U)$ is given by the linear entropy of $\sigma_{U}$ [19, 21].
@@
where $S(\rho||\sigma)={\rm Tr}(\rho\ln\rho-\rho\ln\sigma)$ is the quantum relative entropy. It is noteworthy that the sum of the $\rho_{S}(t)$ and $\rho_{E}(t)$ entropies is not the entropy of the total system as it lacks the entropy contribution contained in the correlations, in particular entanglement between the two subsystems [43]. To account for this, correlation entropy $S_{C}(t)$ is defined. It accounts for the entropy contribution due to entanglement between the system and the reservoir and is given by
@@
For the unitary evolution, the bipartite OTOC, Eq. (5), shows scrambling of information between the two subsystems with the evolution of time as the initial local behavior of the system gets altered.
@@
Figure 9: Comparison of the bipartite OTOC and operator entanglement $E_{op}(U)$ for three different angles, where the bipartition is outlined in Fig. 8, and the total system evolves unitarily. Here, $J=\mathcal{B}=0.5$ and $T=1$.
@@
As can be seen from Fig. 9, the bipartite OTOC increases initially and then varies around a saturation value, except for the case where $\theta=0$ (integrable case). For $\theta=0$, we observe an oscillatory behavior of the bipartite OTOC. The bipartite OTOC is higher for the angle $\theta=7\pi/16$ when compared with the bipartite OTOC for $\theta=\pi/2$. The long-time behavior of the curve for $\theta=\pi/2$ shows deviation from the saturation value, consistent with the behavior observed in Fig. 7. Interestingly, the operator entanglement for all the angles match exactly with the variation of the bipartite OTOC. This benchmarks the statement that the operator entanglement serves as a witness of (non-)integrable dynamics of quantum systems for unitary evolution.
@@
Figure 10: (a) Entropy production $\Sigma$ and (b) correlation entropy $S_{C}$ of the $S$-$E$ Ising spin system for three different angles of tilt of the magnetic field. The parameters are $\mathcal{B}=J=0.5$.
@@
The entropy production and the correlation entropy are plotted in Fig. 10. It is observed that the entropy production $\Sigma$ rises quickly, reminiscent of the behavior of the bipartite OTOC, when the angle of tilt of the magnetic field is non-zero. The highs and the lows of the bipartite OTOC match those of the entropy production. The similarity of bipartite OTOC with entropy production hints that entanglement is the underlying factor that affects the commutativity of any two unitary operators in their respective subspaces and the entropy production. It is noteworthy that the bipartite OTOC reveals the changes through the operator evolution, and the entropy production does it through the state evolution. Interestingly, for some time periods, the entropy production rate is negative, which was used as an identifier of non-Markovianity in [65]; even though in the present case, the evolution considered is unitary.
@@
Figure 11: The quantity $\Sigma+S_{C}$ as a function of time. The parameters are $\mathcal{B}=J=0.5$.
@@
Further, the correlation entropy $S_{C}$ is plotted in Fig. 10(b). It can be seen that the values of $|S_{C}|$ are lesser than the values of the entropy production, that is, $\Sigma+S_{C}\geq 0$. For $\theta=0$, we observe from Figs. 10(a) and (b) that both the entropy production and the correlation entropy are zero, indicating its trivially integrable behavior. For $\theta=\pi/2$, the entropy production and the correlation entropy match exactly, modulo with opposite signs, that is, the sum $\Sigma+S_{C}=0$ (Fig. 11). Hence, the irreversibility due to the lack of non-local information shared between the two subsystems $S$ and $E$ is exactly compensated by the entropy contribution due to entanglement between them. This indicates that the Ising model is non-trivially integrable at the angle $\theta=\pi/2$; in fact, it can be made integrable by mapping it to a model of non-interacting fermions using the Jordan-Wigner transformation. There’s a mismatch between $\Sigma$ and $S_{C}$ for the angle $\theta=7\pi/16$, which is evident from Figs. 10 and 11, indicating a non-integrable behavior of the system. Interestingly, these features of the Ising model are corroborated here by tools of quantum statistical mechanics, that is, entropy production and correlation entropy. We note here that in order to facilitate a comparison with bipartite OTOC, the entropy production and correlation entropy are calculated by taking the $E$ subsystem in the maximally mixed state and the $S$ subsystem in any state, taken here as the ground state. This choice is followed subsequently. The choice of different states affects the difference in values between the bipartite OTOC and the entropy production, even though the qualitative behavior remains similar. Further, the results for $\theta=0$ that both $\Sigma$ and $S_{C}$ are zero and for $\theta=\pi/2$ that the sum $\Sigma+S_{C}=0$ continue to be satisfied for different states.
@@
Figure 12: Variation of the entropy production $\Sigma$, correlation entropy $S_{C}$, and bipartite OTOC $G(\mathcal{U})$ with time for the (a) Dicke and (b) TC models, under unitary evolution. Here, (a) $\lambda=1.5$, and (b) $\lambda=2$, and $\omega_{c}=\omega_{0}=2$.
@@
To further motivate the connection between entropy production and bipartite OTOC, we briefly return to the Dicke and TC models considered above. Figure 12 depicts the entropy production, correlation entropy, and bipartite OTOC for the Dicke and the TC model undergoing unitary evolution. Here, the atomic and the radiation field parts are the subsystems $A$ and $B$, respectively, for the calculation of the bipartite OTOC and are analogous to the symbols $S$ and $E$ for the calculation of the entropy production, Eq. (20), and correlation entropy, Eq. (21). It can be observed that the variations in the entropy production and the bipartite OTOC are in good agreement with each other. This suggests a relationship between irreversibility, characterized by entropy production, and information scrambling, characterized by bipartite OTOC, of the system. However, the numerical values of the two do not coincide, indicating that information scrambling and irreversibility, though similar, are not synonymous. Further, in both cases, the values of the correlation entropy are lesser than the entropy production, and in the case of the Dicke model, $S_{C}\sim-\Sigma$.

### v2

# thanks: Corresponding author

@@ I Introduction
The relationship between the classical and quantum dynamics of chaotic systems is at the heart of the study of quantum chaos. Recently, there has been a resurgence of interest in many-body quantum chaos and its relationship to information scrambling. Information scrambling refers to the spreading of quantum information over a complex quantum system such that it is not accessible to simple probes. Interest in this field was rejuvenated recently by works on black hole information paradox [1, 2], and the problem of quantum thermalization [3]. Traditionally, the ideas of random matrix theory were employed to probe quantum chaos [4, 5]. In recent times, the out-of-time-ordered correlator (OTOC) has emerged as a prominent tool to investigate information scrambling. This was first brought out in the context of superconductivity [6] and was popularized by [7]. The growth rate of the OTOC is commonly attributed to the classical Lyapunov exponent [8]. OTOCs have been extensively used in the study of a number of systems, such as in quantum field theories [9], random unitary models [10], spin chains [11, 12], and quantum optical models [8, 13, 14], among others [15, 16, 17]. Various forms of the OTOC, such as the regularized and the physical OTOC, were also used in the context of the fluctuation-dissipation theorem, generalized to the case of the OTOC [18].
@@ III.1 Dicke model
$\frac{\partial}{\partial t}\mathcal{O}=\mathcal{L}^{\dagger}(\mathcal{O})$ $=i[H_{Dicke},\mathcal{O}]+\gamma(N_{thA}+1)\left[J_{+}\mathcal{O}J_{-}-\frac{1}{2}\left\{\mathcal{O},J_{+}J_{-}\right\}\right]$
$+\gamma N_{thA}\left[J_{-}\mathcal{O}J_{+}-\frac{1}{2}\left\{\mathcal{O},J_{-}J_{+}\right\}\right]$
$+\kappa(N_{thB}+1)\left[a^{\dagger}\mathcal{O}a-\frac{1}{2}\left\{\mathcal{O},a^{\dagger}a\right\}\right]$
$+\kappa N_{thB}\left[a\mathcal{O}a^{\dagger}-\frac{1}{2}\left\{\mathcal{O},aa^{\dagger}\right\}\right],$ (12)
@@
where $\gamma$ acts as a dissipating factor modelling spontaneous emission, and $\kappa$ is responsible for cavity decay. Further, $N_{th_{k}}=\frac{1}{e^{\beta\omega_{k}}-1}$ with $\beta=\frac{1}{k_{B}T}$, and $k=A$, $B$ denoting the subspaces $A$ and $B$, respectively. By vectorizing this master equation, the Lindbladian superoperator $\mathcal{L}^{\dagger}$ and, in turn, the CPTP map $\mathcal{E}^{\dagger}=e^{t\mathcal{L}^{\dagger}}$ is obtained. We consider the atomic system to belong to the $A$ subspace ($\omega_{A}=\omega_{0}$), and the radiation field is relegated to the $B$ subspace ($\omega_{B}=\omega_{C}$). By applying the swap operators for the specific dimensions, the bipartite OTOC is obtained. Further, the Fock space of the field part is truncated in such a manner that increasing the dimension of the Fock space after a point doesn’t affect the dynamics  [13, 62]. The dynamics of the bipartite OTOC for the Dicke model for different values of atom-field interaction strength $\lambda$ is shown in Fig. 2(a).
@@
Figure 2: Variation of averaged bipartite OTOC $G(\mathcal{E}^{\dagger})$ for the Dicke model with time $t$ for (a) different values of spin-cavity field interaction strength $\lambda$, (b) different values of temperature $T$ at $\lambda=0.1$ and (c) at $\lambda=1.3$, and (d) different values of dissipative strengths $\gamma$ and $\kappa$. The value of parameters are: $\omega_{0}=2$, $\omega_{c}=2$.
@@
We observe that as the atom-field coupling strength $\lambda$ increases, the fluctuations (or revivals) in the Bipartite OTOC $G(\mathcal{E}^{\dagger})$ decrease; that is, the curves smooth out. This is predominant when the coupling parameter approaches the critical value $\lambda_{c}=\frac{\sqrt{\omega\omega_{0}}}{2}$, where the Dicke model transitions from the normal regime to the superradiant regime. In [13], an exponential behavior of OTOC was observed when the Dicke model underwent a phase transition from the normal to the superradiant regime. A consistent result is observed here for the bipartite OTOC, where the smoothening of the curve in the superradiant regime, characterized here by $\lambda=1$, could be considered an indicator of the onset of quantum chaos. Further, in Figs. 2(b) and 2(c), we plot the variation of the bipartite OTOC for different temperatures for given values of $\lambda$. It is observed that the bipartite OTOC takes the highest values for zero temperature. As the temperature increases, the values of the bipartite OTOC reduce. Further, in Fig. 2(d), we plot the variation of the bipartite OTOC with time for different values of the dissipative strength $\gamma$ and $\kappa$ at zero temperature. In this case, the bipartite OTOC reduces as we increase the dissipative strength. For $\gamma=\kappa=0$, the bipartite OTOC does not dissipate with time.
@@ III.2 Tavis-Cummings model
The total number of the excitations $\mathcal{N}=a^{\dagger}a+J_{z}$ in this model remains conserved. The sources of dissipation in this model are the same as in the Dicke model. By replacing $H_{Dicke}$ in Eq. (III.1) by $H_{TC}$, we arrive at an adjoint master equation for the TC model. The swap operators for this model remain the same as those for the Dicke model. The variation of the bipartite OTOC for the TC model is similar to the variation of the bipartite OTOC for the Dicke model. Like the Dicke model, as the coupling strength increases, the fluctuations (or revivals) in the bipartite OTOC $G(\mathcal{E}^{\dagger})$ decrease. The fluctuations in the bipartite OTOC act as an identifier of an integrable system; the smoothness of the bipartite OTOC indicates non-integrability.
@@ III.3 Ising model with tilted magnetic field
$H_{A}$ $=\mathcal{B}\left(\sin(\theta)\sigma^{x}_{1}+\cos(\theta)\sigma^{z}_{1}\right),$
$H_{B}$ $=\mathcal{B}\sum_{i=2}^{N}\left(\sin(\theta)\sigma^{x}_{i}+\cos(\theta)\sigma^{z}_{i}\right),~{}~{}~{}~{}\text{and}$
@@
$\frac{\partial}{\partial t}\mathcal{O}$ $=i[H_{\theta},\mathcal{O}]+\sum_{j=1}^{N}\bigg{(}\gamma(N_{th_{j}}+1)\bigg{[}\sigma_{+}^{j}\mathcal{O}\sigma_{-}^{j}-\frac{1}{2}\{\mathcal{O},\sigma_{+}^{j}\sigma_{-}^{j}\}\bigg{]}$
$+\gamma N_{th_{j}}\bigg{[}\sigma_{-}^{j}\mathcal{O}\sigma_{+}^{j}-\frac{1}{2}\{\mathcal{O},\sigma_{-}^{j}\sigma_{+}^{j}\}\bigg{]}\bigg{)},$ (16)
where $N$ is the number of spins in the Ising chain being acted upon by the bath, and $\gamma$ is the dissipation factor. Further, $N_{th_{j}}=\frac{1}{e^{\beta\omega_{j}}-1}$, where $\omega_{j}$ is the transition frequency of the spin at site $j$, which in this case, for each spin is $2\mathcal{B}$.
@@
The variation of the bipartite OTOC for this model for four different angles of tilt, including the two extreme cases of $\theta=0^{\circ}$ and $\theta=90^{\circ}$ is plotted in Fig. 4(a).
Figure 4: Variation of averaged bipartite OTOC for (a) the Ising model in the tilted magnetic field $G(\mathcal{E}^{\dagger})$ with time $t$ for different values of angle of tilt of the magnetic field $\theta$, and (b) a system with two bipartitions with subsystem A having 2 spins and B having 3 spins with the first and last spins experiencing the influence of baths. Two cases are considered in (b), where the two baths are at the same ($T_{1}=T_{5}=1$) or different temperatures ($T_{1}=1$, and $T_{5}=5$). The values of parameters are (a): $T=1$, $\gamma=0.01,\mathcal{B}=J=0.5$, (b): $J=\mathcal{B}=1,\gamma=0.05$.
@@
Further, we study the bipartite OTOC in the above Ising model under the assumption that the bath acts only on the first and the last spins of the Ising chain. The bipartition of the system is taken as 2 spins in the $A$ subsystem and 3 spins in the $B$ subsystem, as illustrated in the second sketch of Fig. 3. Here, we can study a spin chain attached to the baths at the same or different temperatures on either side. In this case, only the first and last spins are subjected to the dissipating Lindblad operators $L_{j}$’s as $\sigma_{\pm}^{j}$ (where $j=1,N)$. To this end, the above master equation [Eq. (III.3)] with appropriate Lindblad operators is used. The bipartite OTOC for this system is shown in Fig. 4(b) for two different angles of tilt with the magnetic field for the same or different temperatures of the two baths. It is evident from the figure that the bipartite OTOC takes lesser values in the presence of different bath temperatures, with the last spin having a higher temperature than the first one. Furthermore, a smooth graph of the bipartite OTOC is obtained for the angle $167\pi/$ 16, indicating non-integrable behavior, whereas the graph for the angle $\pi/2$ has fluctuations in it, indicative of integrability. This is consistent with the (non-)integrable behavior of the Ising model, Eq. (III.3). Interestingly, from the plots of the bipartite OTOC in Figs. 4(a) and (b), for the angle $\theta=\pi/2$, we observe that the detection of integrable behavior is hampered by the dissipation.
@@ IV Bipartite OTOC, operator entanglement, and entropy production
The operator entanglement in many-body quantum systems quantifies how much the local operators become entangled over time. It serves as a natural candidate for distinguishing between chaotic and integrable dynamics in quantum systems [64]. Further, it was shown in [19, 22, 21] that the bipartite OTOC for a unitary evolution $U$ matches the operator entanglement $E_{op}(U)$. The total unitary |$\ket{U}_{ABA^{\prime}B^{\prime}}=U_{AB}\otimes I_{A^{\prime}B^{\prime}}\ket{\psi^{+}}$ for the maximally entangled state |$\ket{\psi^{+}}=\frac{1}{\sqrt{d}}\sum_{j=0}^{d-1}\ket{j}_{AB}\ket{j}_{A^{\prime}B^{\prime}}$ is supported over the composite Hilbert space $\mathcal{H}_{AB}$ and its replica $\mathcal{H}_{A^{\prime}B^{\prime}}$. Since the entanglement of state |$\ket{U}_{ABA^{\prime}B^{\prime}}$ across $AA^{\prime}|BB^{\prime}$ is non-trivial, in contrast to the bipartition $AB|A^{\prime}B^{\prime}$ which is maximal due to it being local unitarily equivalent to the maximally entangled state, $B$ and $B^{\prime}$ are traced out. The explicit expression for $E_{op}(U)$ is given by the linear entropy of $\sigma_{U}$ [19, 21].
@@
where $S\left[(*)||(\circ)\right]={\rm Tr}\left[(*)\ln(*)-(*)\ln(\circ)\right]$ is the quantum relative entropy. It is noteworthy that the sum of the $\rho_{S}(t)$ and $\rho_{E}(t)$ entropies is not the entropy of the total system as it lacks the entropy contribution contained in the correlations, in particular entanglement between the two subsystems [43]. To account for this, correlation entropy $S_{C}(t)$ is defined. It accounts for the entropy contribution due to entanglement between the system and the reservoir and is given by
@@
For the unitary evolution, the bipartite OTOC, Eq. (5), shows the scrambling of information between the two subsystems with the evolution of time as the initial local behavior of the system gets altered.
@@
Figure 6: Comparison of the bipartite OTOC and operator entanglement $E_{op}(U)$ for three different angles, where the bipartition is outlined in Fig. 5, and the total system evolves unitarily. Here, $J=\mathcal{B}=0.5$.
@@
As can be seen from Fig. 6, the bipartite OTOC increases initially and then varies around a saturation value, except for the case where $\theta=0$ (integrable case). For $\theta=0$, we observe an oscillatory behavior of the bipartite OTOC. The bipartite OTOC is higher for the angle $\theta=7\pi/16$ when compared with the bipartite OTOC for $\theta=\pi/2$. The long-time behavior of the curve for $\theta=\pi/2$ shows deviation from the saturation value, consistent with the behavior observed in Fig. 4. Interestingly, the operator entanglement for all the angles matches exactly with the variation of the bipartite OTOC. This benchmarks the statement that the operator entanglement serves as a witness of (non-)integrable dynamics of quantum systems for unitary evolution.
Figure 7: The bipartite OTOC for unitary evolution of a system of 4 spins having two different bipartitions: $A:B=1:3$ and $2:2$. In (a), we take the case for the angle of tilt $\theta=0$ (trivially integrable) and $\theta=\pi/2$ (non-trivially integrable). In (b), we take two angles, $\theta=\pi/4$ and $\theta=7\pi/16$ (non-integrable). The parameters are taken as $J=\mathcal{B}=0.5$.
In Fig. 7, we compare the bipartite OTOC for two different bipartitions $A:B=1:3$ and $2:2$ for the system of 4 spins, Fig. 5. We observe that the bipartite OTOC always takes higher values for the 2:2 bipartition case as compared to the 1:3 bipartition case. Further, the qualitative behavior of the bipartite OTOC is similar for $\theta=0$ and $\pi/2$ for both kinds of bipartitions. This behavior is absent in the non-integrable regime.
@@
Figure 8: (a) Entropy production $\Sigma$ along with bipartite OTOC $G(t)$ and (b) correlation entropy $S_{C}$ of the $S$-$E$ Ising spin system for three different angles of tilt of the magnetic field. The parameters are $\mathcal{B}=J=0.5$.
@@
The entropy production $\Sigma$ and the correlation entropy $S_{C}$ for the tilted Ising model are calculated using the ground state as the initial state for the subsystem $S$ and the maximally mixed state for the subsystem $E$. Both are plotted in Fig. 8 along with the bipartite OTOC $G(t)$. It is observed that the entropy production $\Sigma$ rises quickly, reminiscent of the behavior of the bipartite OTOC, when the angle of tilt of the magnetic field is non-zero. The highs and the lows of the bipartite OTOC match those of the entropy production. The similarity of bipartite OTOC with entropy production hints that entanglement is the underlying factor that affects the commutativity of any two unitary operators in their respective subspaces and the entropy production. It is noteworthy that the bipartite OTOC reveals the changes through the operator evolution, and the entropy production does it through the state evolution. This is consistent with the observation made in [65], where a connection between the bipartite OTOC and the mutual information between the subsystems was established. It can be pointed out here that the entropy production, Eq. (20), can also be written as the sum of the mutual information between the subsystems $S$ and $E$ and the Kullback Leibler divergence between the states $\rho_{E}(t)$ and $\rho_{E}(0)$. Therefore, analogous to the relationship between the bipartite OTOC and the mutual information, a similar connection between the entropy production and the bipartite OTOC can be established. Interestingly, for some time periods, the entropy production rate is negative, which was used as an identifier of non-Markovianity in [66].
@@
Figure 9: The quantity $\Sigma+S_{C}$ as a function of time for the tilted field Ising model. The parameters are $\mathcal{B}=J=0.5$.
@@
Further, the correlation entropy $S_{C}$ is plotted in Fig. 8(b). It can be seen that the values of $|S_{C}|$ are lesser than the values of the entropy production, that is, $\Sigma+S_{C}\geq 0$. For $\theta=0$, we observe from Figs. 8(a) and (b) that both the entropy production and the correlation entropy are zero, indicating its trivially integrable behavior. For $\theta=\pi/2$, the entropy production and the correlation entropy match exactly, modulo with opposite signs, that is, the sum $\Sigma+S_{C}=0$ (Fig. 9). Hence, the irreversibility due to the lack of non-local information shared between the two subsystems $S$ and $E$ is exactly compensated by the entropy contribution due to entanglement between them. This indicates that the Ising model is non-trivially integrable at the angle $\theta=\pi/2$; in fact, it can be made integrable by mapping it to a model of non-interacting fermions using the Jordan-Wigner transformation. There’s a mismatch between $\Sigma$ and $S_{C}$ for the angle $\theta=7\pi/16$, which is evident from Figs. 8 and 9, indicating a non-integrable behavior of the system. Interestingly, these features of the Ising model are corroborated here by tools of quantum statistical mechanics, that is, entropy production and correlation entropy. We note here that in order to facilitate a comparison with bipartite OTOC, the entropy production and correlation entropy are calculated by taking the $E$ subsystem in the maximally mixed state and the $S$ subsystem in any state, taken here as the ground state. This choice is followed subsequently. The choice of different states affects the difference in values between the bipartite OTOC and the entropy production, even though the qualitative behavior remains similar.
Figure 10: Variation of the entropy production for the tilted field Ising model, Fig. 5, in the weak $S$ and $E$ coupling scenario with the number of spins increased to 6 in the $E$ subsystems. The values of $\mathcal{B}=J=0.5$ and $\theta=\pi/2$. Further, interaction strength between the spins 1 and 2 is taken to be $10^{-3}J$.
Also, we plot the variation of the entropy production and the correlation entropy for the tilted field Ising model with weak $S$ and $E$ coupling strength for the angle $\theta=\pi/2$ in Fig. 10. This is done to simulate the impact of the GKSL type of evolution on the entropy production. It is observed that the values of entropy production are very low (of the order of $10^{-5}$) as compared to the previous cases, and they increase as time progresses. Again, in this case of $\theta=\pi/2$, the sum of the entropy production and the correlation entropy is zero.
@@
Figure 11: Variation of the entropy production $\Sigma$, correlation entropy $S_{C}$, and bipartite OTOC $G(t)$ with time for the (a) Dicke and (b) TC models, under unitary evolution. Here, (a) $\lambda=1.5$, and (b) $\lambda=2$, and $\omega_{c}=\omega_{0}=2$.
@@
To further motivate the connection between entropy production and bipartite OTOC, we briefly return to the Dicke and TC models considered above. Figure 11 depicts the entropy production, correlation entropy, and bipartite OTOC for the Dicke and the TC model undergoing unitary evolution. Here, the atomic and the radiation field parts are the subsystems $A$ and $B$, respectively, for the calculation of the bipartite OTOC and are analogous to the symbols $S$ and $E$ for the calculation of the entropy production, Eq. (20), and correlation entropy, Eq. (21). Here, the subsystems $S$ and $E$ are initially taken to be in the ground state and the maximally mixed state, respectively. It can be observed that the variations in the entropy production and the bipartite OTOC are in good agreement with each other. This suggests a relationship between irreversibility, characterized by entropy production, and information scrambling, characterized by bipartite OTOC, of the system. However, the numerical values of the two do not coincide, indicating that information scrambling and irreversibility, though similar, are not synonymous. Further, in both cases, the values of the correlation entropy are lesser than the entropy production, and in the case of the Dicke model, $S_{C}\sim-\Sigma$.